# Label Leakage in Federated Inertial-based Human Activity Recognition

### v1

@@ Abstract.
While prior work has shown that Federated Learning updates can leak sensitive information, label reconstruction attacks, which aim to recover input labels from shared gradients, have not yet been examined in the context of Human Activity Recognition (HAR). Given the sensitive nature of activity labels, this study evaluates the effectiveness of state-of-the-art gradient-based label leakage attacks on HAR benchmark datasets. Our findings show that the number of activity classes, sampling strategy, and class imbalance are critical factors influencing the extent of label leakage, with reconstruction accuracies reaching up to 90% on two benchmark datasets, even for trained models. Moreover, we find that Local Differential Privacy techniques such as gradient noise and clipping offer only limited protection, as certain attacks still reliably infer both majority and minority class labels. We conclude by offering practical recommendations for the privacy-aware deployment of federated HAR systems and identify open challenges for future research. Code to reproduce our experiments is publicly available via github.com/mariusbock/leakage_har.
@@ 1. Introduction
We demonstrate that the unbalanced nature of HAR datasets introduces unique privacy risks in federated settings, and we assess how this necessitates stronger applications of Local Differential Privacy (LDP) methods to effectively hide user label information.
@@ 3. Methodology
We first formulate the problem setting and discuss the possible attacks we consider in this area, before we focus on the defense measures and the experimental setup of our study.
@@ 3.1. Problem Setting & Threat Model
based on local training data , a suitable loss function , and performs one or multiple steps of gradient descent on . The new parameters are subsequently sent to the server. In the simplest case of a single gradient descent step, the data sent to the server are the gradients of the costs w.r.t. the weights and biases of the network. In this work, we study what this information reveals about the empirical distribution of the activity labels in the client’s training data . Specifically, we focus on the reconstruction of label information from gradients of the last layer through gradient inversion attacks, as detailed in Section 3.2. Throughout our work, we assume that clients compute and share correct gradients without using class weights based on valid data and labels and also assume that the server knows the number of training examples , which were used to compute the individual gradients. The threat model considered is that of an honest-but-curious server: the server performs protocol-compliant aggregation but is passively interested in extracting sensitive information from the received updates (see Figure 1). While alternative threat models, such as an actively malicious server that modifies the model to enhance information leakage, or clients that train with more than data points, are plausible, they fall outside the scope of this work.
@@ 3.3. Defense Measures
Though research has continuously worked on Differential Privacy (DP) methods (McMahan et al., 2018) which are less likely to be compromised, such as secure aggregation (Bonawitz et al., 2019), many of these methods require coordination of the server, i.e. requiring trust of the users in the server. The Federated Learning environment sketched in Figure 1 assumes that users do not trust the server they are sending their gradient-based updates to. We thus only focus on Local Differential Privacy methods, namely gradient clipping and noise addition. These methods, being executed locally by each user individually, do not require users to trust the server. In a realistic setting, a privacy budget determines the degree of LDP to be applied during the federated training process of the neural network. This work, though, applies the approach of Gat et al. (Gat and Sharif, 2024) where no privacy budget needs to be retained because the label reconstruction is evaluated on single batches sampled from the datasets without running a complete training scenario. The clipping of the gradients is done by normalizing the gradient vectors to an L2-norm of if the norm of the gradient exceeds a defined threshold. Noise is added to the gradients by sampling from a Gaussian distribution that has a mean of and a predefined standard deviation of .
@@ 3.4. Experimental Setup
and average these results over all of a client to compute a single LeAcc and LeAcc metric. As LeAcc and LnAcc are compute per batch, datasets which show an unbalanced class count can cause these metrics to be biased towards majority classes. We thus propose a third metric, Class-wise Average Accuracy (ClassAcc), to evaluate leakage results. Unlike LeAcc and LnAcc, ClassAcc is computed on the complete dataset of a client, i.e. all that were reconstructed of a specific client. The metric is computed by calculating the class-wise accuracy between the label-sorted, concatenated reconstructed labels with the ground truth labels of all .
@@ 4. Results
Comparison of LnAcc, LeAcc and ClassAcc of the investigated attacks (Gat and Sharif, 2024; Wainakh et al., 2021; Ma et al., 2023) using a shuffling sampling technique and a batch size of 100 samples. Each value is the average across the different dataset (WEAR (Bock et al., 2024a) and Wetlab (Scholl et al., 2015)) and architecture combinations. One can see that balanced sampling negatively influences the LnAcc of a majority of the label leakage attacks.
@@
Figure 2 presents the results of the evaluated label leakage attacks on the WEAR and Wetlab datasets, using shuffled sampling with a batch size of 100 sliding windows. Overall, bias-based techniques, namely EBI, iLRG, and LLBG, consistently outperform the weight-based techniques LLG and LLG* by a significant margin. Among all methods, iLRG achieves the highest performance on untrained models, with an average ClassAcc of approximately 90% on the WEAR dataset and 80% on the Wetlab dataset. In contrast, for trained models, EBI and LLBG are the most effective attacks, achieving around 55% on WEAR and 50% on Wetlab. In addition to the shuffled sampling approach, we evaluate two alternative sampling techniques: sequential and balanced. Sequential sampling preserves the temporal order of sliding windows, while balanced sampling constructs batches containing an equal number of samples per class by randomly drawing from class-specific pools. Note that to address class imbalance in the dataset, balanced batches are filled using repeated samples as needed.
Table 1 reports the average performance for each sampling technique, aggregated across bias- and weight-based attacks and model architectures. Results show that sequentially sampled batches exhibit significantly more label leakage than balanced or shuffled batches. When comparing the performance of attacks on untrained versus trained models, LLBG and EBI show improved performance on trained models when using shuffled and balanced sampling. In contrast, techniques such as iLRG experience a noticeable drop in effectiveness when applied to trained models. Among all methods, EBI remains the most stable across both trained and untrained scenarios. When comparing the two architectures, TinyHAR and DeepConvLSTM, performance varies across experiments. However, TinyHAR consistently shows greater vulnerability to the iLRG attack, with higher leakage levels across all evaluations. Meanwhile, EBI and LLBG generally perform better on the DeepConvLSTM than on TinyHAR when models are trained. Looking at dataset differences, the WEAR dataset consistently exhibits greater label leakage than Wetlab under sequential sampling. In particular, EBI and LLBG maintain ClassAcc levels above 90% on WEAR, even with trained models. This heightened vulnerability is likely due to the recording setup of the WEAR dataset, which contains fewer activity transitions. As a result, sequential batches more frequently become label-exclusive, i.e., containing only a single class. Our results indicate that such batches pose a serious privacy risk, as labels can be accurately reconstructed in various settings. In summary, since shuffled batches consistently yield the lowest leakage, we hypothesize that effective mitigation of label leakage in HAR involves constructing batches that contain a diverse set of activity classes, yet are not fully class-balanced.
@@
Shuffle | Weight | 7.11% | 7.13% | 14.64% | 15.72% |
Bias | 32.79% | 15.51% | 28.20% | 27.87% | |
Avg. | 19.95% | 11.32% | 21.42% | 21.66% | |
Sequential | Weight | 26.79% | 39.76% | 23.94% | 28.88% |
Bias | 98.50% | 77.54% | 79.00% | 66.75% | |
Avg. | 62.65% | 58.65% | 51.47% | 47.82% | |
Balanced | Weight | 7.00% | 27.66% | 12.83% | 13.09% |
Bias | 48.93% | 54.24% | 52.24% | 44.50% | |
Avg. | 27.97% | 40.95% | 32.53% | 28.80% |
Exemplary multi-step experimental results applied on the Wetlab dataset for an untrained DeepConvLSTM. We report the ClassAcc difference of the investigated attacks with the single-step experiment leakage using a batch size () of 100. We test splitting said batch size into 2 and 5 update steps () as well as a larger batch size (500) with 5 update steps. Biased-based attacks, except iLRG, remain largely unaffected by multi-step updates. In case of sequential sampling, label leakage can be significantly reduced when using a larger as batches become less class-exclusive.
@@
Since communication costs between clients and server are a critical resource in federated learning, McMahan et al. (McMahan et al., 2017) proposed that clients perform multiple local gradient update steps before communicating with the server. This approach, known as FedAVG, significantly reduces communication overhead, as clients no longer need to transmit updates after every local step. In the context of our investigation into label leakage attacks, we assume that the client and server agree on a fixed number of data points, , for local training before sending an update. We explore whether averaging gradients over local updates, i.e., splitting into equal-sized mini-batches, can reduce label leakage risks from a single user.
Figure 3 presents sample results from our multi-step experiments using the Wetlab dataset and an untrained DeepConvLSTM model. The results show that bias-based attacks, with the exception of iLRG, maintain stable performance even when reconstructing multi-step updates from untrained models. When applied to trained models however, LLBG and EBI exhibit more significant reductions in leakage, particularly under balanced sampling. Notably, increasing results in the largest performance drop in ClassAcc and LnAcc for sequential sampling, which we hypothesize is due to batches becoming less class-exclusive as more data points are included, thereby reducing the effectiveness of reconstruction. Interestingly, iLRG, while heavily affected by multi-step updates, proves to be the more stable when applied to a trained models, and even increases in performance in certain cases when applied to shuffled and balanced batches. Finally, our results indicate that single-step and multi-step updates pose similar privacy risks. Though larger amounts of local client data showed to decrease leakage, it does not hide users label information effectively with LnAcc and LeAcc still remaining high, suggesting certain classes still being able to be reconstructed.
Results of the LLBG label leakage attack (Gat and Sharif, 2024) being applied to the Wetlab dataset (Scholl et al., 2015) using a trained DeepConvLSTM (Ordóñez and Roggen, 2016). Confusion matrices show the reconstructed label accuracy for the shuffled and sequential sampling case when using (1) no local differential privacy, (2) gaussian noise addition (3) gradient clipping and (4) the combination of both as described in the paper. One can see that the combination of noise and clipping can provide substantial label privacy in both sampling scenarios, yet does not provide full label privacy, with certain sampling methods still showing vulnerabilities.
@@
Having shown that multi-step updates alone do not sufficiently mitigate label leakage, we further investigate whether LDP techniques, specifically gradient clipping and gradient noise, can better protect clients’ label information from the server. We repeated our initial experiments, this time applying one of the following LDP configurations: (1) gradient noise (Gaussian noise with mean 0 and standard deviation 0.1), (2) gradient clipping (scaling gradient vectors to an L2-norm of 0.1), or (3) a combination of both. Importantly, LDP measures were applied only to the gradients of the last model layer, as these are targeted by the label leakage attacks. Figure 4 presents representative results for the LLBG attack applied to a trained DeepConvLSTM model. Overall, when comparing across all attack methods, we observe that gradient clipping alone is largely ineffective, particularly on gradients from sequentially sampled batches. With the exception of iLRG, which consistently shows unstable performance and is heavily impacted by both clipping and noise, most other methods exhibit minimal change in effectiveness with clipping applied. Among all techniques, EBI remains the most stable, maintaining consistent performance across all LDP variants.
Sequential sampling continues to be the most vulnerable configuration, with EBI and LLBG achieving around 70% ClassAcc for untrained models and approximately 60% for trained models, even in the presence of noise. Notably, only the combination of clipping and noise proves to be an effective defense under sequential sampling. In contrast, noise addition alone significantly degrades performance under balanced and shuffled sampling, with ClassAcc dropping below 30% and nearing 20%, respectively. Although noise reduces LnAcc, suggesting less precise label prediction, LeAcc remains high, indicating that attackers can still reliably infer whether a label is present in the batch. Even though the combined use of clipping and noise provides a meaningful defense against bias-based attacks on sequential batches, LeAcc scores remain above 50% for all leakage attacks revealing vulnerabilities that the presence of classes can still be recovered with higher accuracy, raising ongoing privacy concerns despite the use of LDP. This further underscores the need for more robust privacy-preserving strategies, especially for protecting less frequent activity classes.
@@ 5. Limitations
Our paper presents the first comprehensive benchmarking of popular label leakage attacks on HAR benchmark datasets. While the effectiveness of these methods varies across architectures and datasets, our findings reveal a consistent trend: the unique characteristics of HAR data introduce significant privacy vulnerabilities that warrant serious attention. In particular, the inherent class imbalance in HAR datasets limits the reliability of conventional metrics such as LnAcc and LeAcc, commonly used in computer vision. These metrics tend to favor majority classes and can thus give a misleading impression of attack effectiveness. To address this, we introduced ClassAcc, a class-averaged metric that is calculated on a dataset-level, which contributes to a more class-balanced view of exhibited label leakage. Our evaluation of LDP defenses showed that while gradient noise and clipping are limited in isolation, their combination offers a more effective defense. Although more advanced privacy-preserving mechanisms exist, we chose these techniques due to their widespread adoption and computational simplicity. Additional experiments (not detailed in the main paper) further indicate that applying clipping and noise to the gradients of the last layer does not impair the prediction accuracy of trained models. Nevertheless, certain attacks still achieved high recognition rates of class presence. We therefore leave it to future work to explore more robust LDP strategies that can ensure privacy protection in these corner cases.
As discussed in Chapter 3.1, our experiments focus on single-client gradient updates. Future research could examine how multi-client aggregation impacts label leakage, particularly in real-world FL deployments. Additionally, since bias-based attacks consistently outperformed weight-based ones, a practical defense recommendation may be to train HAR models without bias terms, as suggested in (Scheliga et al., 2022). Moreover, although commonly employed in HAR tasks, we deliberately did not apply a weighted loss during training. Weighted losses increase the influence of rare classes on the gradient, potentially making these classes more susceptible to leakage, an especially critical concern in sensitive domains such as healthcare and disease detection. Lastly, given the scope of our experimental framework, we report only the most representative trends in the paper. To support reproducibility and future research, we have open-sourced all code and experiments.

### v2

@@ Abstract.
While prior work has shown that Federated Learning updates can leak sensitive information, label reconstruction attacks, which aim to recover input labels from shared gradients, have not yet been examined in the context of Human Activity Recognition (HAR). Given the sensitive nature of activity labels, this study evaluates the effectiveness of state-of-the-art gradient-based label leakage attacks on HAR benchmark datasets. Our findings show that the number of activity classes, sampling strategy, and class imbalance are critical factors influencing the extent of label leakage, with reconstruction accuracies reaching well-above 90% on two benchmark datasets, even for trained models. Moreover, we find that Local Differential Privacy techniques such as gradient noise and clipping offer only limited protection, as certain attacks still reliably infer both majority and minority class labels. We conclude by offering practical recommendations for the privacy-aware deployment of federated HAR systems and identify open challenges for future research. Code to reproduce our experiments is publicly available via github.com/mariusbock/leakage_har.
@@ 1. Introduction
We show that the unbalanced nature of HAR datasets introduces unique privacy risks in federated settings, and we assess how this necessitates stronger applications of Local Differential Privacy (LDP) methods to effectively hide user label information.
@@ 3.1. Problem Setting & Threat Model
based on local client training data , a suitable loss function , and performs one or multiple steps of gradient descent on . The new parameters are subsequently sent to the server. In the simplest case of a single gradient descent step, the data sent to the server are the gradients of the costs w.r.t. the weights and biases of the network. In this work, we study what this information reveals about the empirical distribution of the activity labels in the client’s training data . Specifically, we focus on the reconstruction of label information from gradients of the last layer through gradient inversion attacks, as detailed in Section 3.2. Throughout our work, we assume that clients compute and share correct gradients without using class weights based on valid data and labels and also assume that the server knows the number of training examples , which were used to compute the individual gradients. The threat model considered is that of an honest-but-curious server: the server performs protocol-compliant aggregation but is passively interested in extracting sensitive information from the received updates (see Figure 1). While alternative threat models, such as an actively malicious server that modifies the model to enhance information leakage, or clients that train with more than data points, are plausible, they fall outside the scope of this work.
@@ 3.3. Defense Measures
Though research has continuously worked on privacy-enhancing methods such as Differential Privacy (DP) (McMahan et al., 2018) and encryption-based techniques (Bonawitz et al., 2019), many of these methods require coordination of the server, i.e. requiring trust of the users in the server. The Federated Learning environment sketched in Figure 1 assumes that users do not trust the server they are sending their gradient-based updates to. We thus only focus on Local Differential Privacy methods, namely gradient clipping and noise addition. These methods, being executed locally by each user individually, do not require users to trust the server. In a realistic setting, a privacy budget determines the degree of LDP to be applied during the federated training process of the neural network. This work, though, applies the approach of Gat et al. (Gat and Sharif, 2024) where no privacy budget needs to be retained because the label reconstruction is evaluated on single batches sampled from the datasets without running a complete training scenario. The clipping of the gradients is done by normalizing the gradient vectors to an L2-norm of if the norm of the gradient exceeds a defined threshold. Noise is added to the gradients by sampling from a Gaussian distribution that has a mean of and a predefined standard deviation of .
@@ 3.4. Experimental Setup
and average these results over all of a client to compute a single LeAcc and LeAcc metric. As LeAcc and LnAcc are compute per batch, datasets which show an unbalanced class count can cause these metrics to be biased towards majority classes. We thus propose a third metric, Class-wise Average Accuracy (ClassAcc), which is calculated as the average of the per-class LnAcc computed across the complete dataset of a client, i.e. all that were reconstructed of a specific client.
@@ 4. Results
Comparison of LnAcc, LeAcc and ClassAcc of the investigated attacks using a shuffling sampling technique and a batch size of 100 samples. Each value is the average across the different dataset (WEAR and Wetlab) and architecture combinations. Error bars refer to Standard Deviation (SD) across clients.
@@
Figure 2 presents the results of the evaluated label leakage attacks on the WEAR and Wetlab datasets, using shuffled sampling with a batch size of 100 sliding windows. Overall, bias-based techniques, namely EBI, iLRG, and LLBG, consistently outperform the weight-based techniques LLG and LLG* by a significant margin. Among all methods, iLRG achieves the highest performance on untrained models, with an average ClassAcc of approximately 97% on the WEAR dataset and 81% on the Wetlab dataset. In contrast, for trained models, EBI and LLBG also become effective attacks, achieving around 70% on WEAR and between 40-47% on Wetlab. In addition to the shuffled sampling approach, we evaluate two alternative sampling techniques: sequential and balanced. Sequential sampling preserves the temporal order of sliding windows, while balanced sampling constructs batches containing an equal number of samples per class by randomly drawing from class-specific pools. Note that to address class imbalance in the dataset, balanced batches are filled using repeated samples as needed.
Table 1 reports the average performance for each sampling technique, aggregated across bias- and weight-based attacks and model architectures. Results show that sequentially sampled batches exhibit significantly more label leakage than balanced or shuffled batches. When comparing the performance of attacks on untrained versus trained models, LLBG and EBI show improved performance on trained models when using shuffled and balanced sampling. In contrast, iLRG experiences a noticeable drop in effectiveness when applied to trained models. Among all methods, EBI remains the most stable across both trained and untrained scenarios. When comparing the two architectures, TinyHAR and DeepConvLSTM, performance varies across experiments. However, a trained TinyHAR model shows greater vulnerability to the iLRG attack, with higher leakage levels across all evaluations. Meanwhile, EBI and LLBG generally perform better on the DeepConvLSTM than on TinyHAR when models are trained. Looking at dataset differences, the WEAR dataset consistently exhibits greater label leakage than Wetlab. In particular under sequential sampling, EBI and LLBG maintain ClassAcc levels above 90% on WEAR, even with trained models. This heightened vulnerability is likely due to the recording setup of the WEAR dataset, which contains fewer activity transitions. As a result, sequential batches more frequently become label-exclusive, i.e., containing only a single class. Our results indicate that such batches pose a serious privacy risk, as labels can be accurately reconstructed in various settings. In summary, since shuffled batches consistently yield the lowest leakage, we hypothesize that effective mitigation of label leakage in HAR involves constructing batches that contain a diverse set of activity classes, yet are not fully class-balanced.
@@
| Shuffle | Weight | 19.09% | 19.13% | 25.89% | 27.12% |
| Bias | 48.50% | 71.21% | 37.76% | 46.67% | |
| Avg. | 36.74% | 50.38% | 33.01% | 38.85% | |
| Sequential | Weight | 27.26% | 40.32% | 25.05% | 30.75% |
| Bias | 98.64% | 80.71% | 81.67% | 70.76% | |
| Avg. | 70.09% | 64.56% | 59.02% | 54.75% | |
| Balanced | Weight | 14.34% | 14.98% | 15.65% | 15.91% |
| Bias | 70.07% | 71.00% | 68.33% | 65.94% | |
| Avg. | 47.77% | 48.59% | 47.26% | 45.93% | |
Exemplary multi-step experimental results applied on the Wetlab dataset for an untrained DeepConvLSTM. We report the average LnAcc difference between single step and multi-step gradient updates using different amounts of client data () and local update steps () for multi-step and fixing single-step experiments at and . Positive values indicate increased leakage in multi-step updates, while negative values indicate reduced leakage.
@@
Since communication costs between clients and server are a critical resource in federated learning, McMahan et al. (McMahan et al., 2017) proposed that clients perform multiple local gradient update steps before communicating with the server. This approach, known as FedAVG, significantly reduces communication overhead, as clients no longer need to transmit updates after every local step. In the context of our investigation into label leakage attacks, we again assume that the client and server agree on a fixed number of data points, , for local training before sending an update. We explore whether averaging gradients over local updates, i.e., splitting into equal-sized mini-batches, can reduce label leakage risks from a single user.
Figure 3 shows sample results comparing the difference in performance between single-step gradient update experiments () with multi-step variants. Results show that bias-based attacks, with the exception of iLRG, maintain stable performance even when reconstructing multi-step updates from untrained models. When applied to trained models however, LLBG and EBI exhibit more significant reductions in leakage, particularly under balanced sampling. Notably, an increased results in the largest performance drop in ClassAcc and LnAcc for sequential sampling, which we hypothesize is due to batches becoming less class-exclusive as more data points are included, thereby reducing the effectiveness of reconstruction. Interestingly, iLRG, while heavily affected by multi-step updates, proves to be the more stable when applied to a trained models, and even increases in performance in certain cases when applied to shuffled and balanced batches. Finally, our results indicate that single-step and multi-step updates pose similar privacy risks. Though larger amounts of local client data showed to decrease leakage, it does not hide users label information effectively with LnAcc and LeAcc still remaining high, suggesting certain classes still being able to be reconstructed.
Results of the LLBG label leakage attack being applied to the Wetlab dataset using a trained DeepConvLSTM. Confusion matrices show the reconstructed label accuracy for the shuffled and sequential sampling case when using (1) no local differential privacy, (2) gaussian noise addition (3) gradient clipping and (4) the combination of both as described in the paper.
@@
Having shown that multi-step updates alone do not sufficiently mitigate label leakage, we further investigate whether LDP techniques, specifically gradient clipping and gradient noise, can better protect clients’ label information from the server. We repeated our initial experiments, this time applying one of the following LDP configurations: (1) gradient noise (Gaussian noise with mean 0 and standard deviation 0.1), (2) gradient clipping (scaling gradient vectors to an L2-norm of 0.1), or (3) a combination of both. Importantly, LDP measures were applied only to the gradients of the last model layer, as these are targeted by the label leakage attacks. Figure 4 presents representative results for the LLBG attack applied to a trained DeepConvLSTM model. Overall, when comparing across all attack methods, we observe that gradient clipping alone is largely ineffective, particularly on gradients from sequentially sampled batches. With the exception of iLRG, other methods exhibit minimal change in effectiveness with clipping applied. Among all techniques, EBI remains the most stable, maintaining consistent performance across all LDP variants.
Sequential sampling continues to be the most vulnerable configuration, with EBI and LLBG achieving more than 70% ClassAcc for untrained models and more than 60% for trained models, even in the presence of noise. Nevertheless, noise addition significantly reduces LnAcc across all sampling methods, suggesting less precise label prediction. However, LeAcc remains high, indicating that attackers can still reliably infer whether a label is present in the batch. Notably, only the combination of clipping and noise proves to be an effective defense under sequential sampling. While this combined approach offers meaningful defense against bias-based attacks on sequential batches, LeAcc scores still exceed 50% across all leakage attacks and even up to 80% in case of the iLRG attack on the Wetlab dataset. This indicates that the presence of classes can still be recovered with relatively high accuracy, raising persistent privacy concerns despite the application of LDP and underscoring the need for more robust privacy-preserving strategies, especially for protecting less frequent activity classes.
@@ 5. Limitations
Our paper presents the first comprehensive benchmarking of popular label leakage attacks on HAR benchmark datasets. While the effectiveness of these methods varies across architectures and datasets, our findings reveal a consistent trend: the unique characteristics of HAR data introduce significant privacy vulnerabilities that warrant serious attention. In particular, the inherent class imbalance in HAR datasets limits the reliability of conventional metrics such as LnAcc and LeAcc, commonly used in computer vision. These metrics tend to favor majority classes and can thus give a misleading impression of attack effectiveness. To address this, we introduced ClassAcc, a class-averaged metric that is calculated on a dataset-level, which contributes to a more class-balanced view of exhibited label leakage. Our evaluation of LDP defenses showed that while gradient noise and clipping are limited in isolation, their combination offers a more effective defense, yet certain attacks still achieved high recognition rates of class presence. Although more advanced privacy-preserving mechanisms exist, we chose these techniques due to their widespread adoption and computational simplicity and leave it to future work to explore more robust LDP strategies that can ensure privacy protection also in case of class presence.
Additional experiments (provided in the code repository) further indicate that applying clipping and noise to the gradients of the last layer does not impair the prediction accuracy of trained models. As our experiments focus on single-client gradient updates, future research could examine how multi-client aggregation impacts label leakage, particularly in real-world FL deployments. Additionally, since bias-based attacks consistently outperformed weight-based ones, a practical defense recommendation may be to train HAR models without bias terms, as suggested in (Scheliga et al., 2022). Further investigation could also be warranted into why iLRG was the most volatile among all label leakage attacks studied, with its performance highly sensitive to LDP measures and multi-step averaging. Moreover, although commonly employed in HAR tasks, we deliberately did not apply a weighted loss during training. Weighted losses increase the influence of rare classes on the gradient, potentially making these classes more susceptible to leakage, an especially critical concern in sensitive domains such as healthcare and disease detection. Lastly, given the scope of our experimental framework, we report only the most representative trends in the paper. To support reproducibility and future research, we have open-sourced all code and experiments in our code repository.